Two bodies of masses 2 kg and 3 kg are connected by a metal wire of cross section 0.04 mm2 . Breaking stress of metal wire is 2.5 GPa . Find the maximum force that can be applied to the mass 3 kg so that the string does not break.

Two bodies of masses 2 kg and 3 kg are connected by a metal wire of cross section 0.04 mm2Breaking stress of metal wire is 2.5 GPa. Find the maximum force that can be applied to the mass 3 kg so that the string does not break.

Grade:8

2 Answers

npcyh
7 years ago
let F be the force applied on 3 kg mass. then F-T=3a (T is the tension in the string and a is acceleration of blocks). and T=2a i.e., a=T/2. putting in above equation F-T=3T/2. i.e., F=5T/2. now maximum stress is T/A=2.5*10^9. so,T=(2.5*10^9)(0.04*10^-6)=100N. so, maximum force=5*100/2=250N. see if i have done any calculation mistake but this is the process
eswar
7 years ago
Two bodies of masses 1 kg and 2 kg are connected by a metal wire shown in figure. A force of 10 N is applied on the body of mass 2 kg. the breaking stress of metal wire is 2 \times {10^9}N{m^{ - 2}}2−mN901×22×109Nm2what should be minimum radius of the wire used, if it is not to break?